OpenStudy (anonymous):

Susan has $4 in nickels and quarters. She has 4 more quarters than nickels. How many of each does she have?

OpenStudy (anonymous):

i love ur name

OpenStudy (anonymous):

system of equations .05n + .25q = 4 q = n + 4 Okay with this?

OpenStudy (mathmate):

For checking the answer without pen and paper: take out the 4 extra quarters, or $4-4*0.25=$3 Now quarters and nickels are in equal numbers to make $3, or $3/(0.25+0.05)=10 of each. So 10 nickels and 14 quarters.